# CPsyCoun: A Report-based Multi-turn Dialogue Reconstruction and Evaluation Framework for Chinese Psychological Counseling

### v2

@@ Abstract
Chenhao Zhang1,3thanks: Equal contribution.thanks: Work done on the Science and Technology Innovation Project of UCAS directed by SIAT.and Renhao Li2,311footnotemark: 1and Minghuan Tan3thanks: Corresponding author.and Min Yang [^3] footnotemark: 3and Jingwei Zhu4and Di Yang4and Jiahao Zhao3,522footnotemark: 2and Guancheng Ye622footnotemark: 2and Chengming Li7and Xiping Hu7and Derek F. Wong2 1 Huazhong University of Science and Technology 2 University of Macau 3 Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 4 University of Science and Technology of China 5 Jilin University 6 South China University of Technology 7 Shenzhen MSU-BIT University ch_zhang@hust.edu.cn, li.renhao@connect.um.edu.mo {mh.tan,min.yang}@siat.ac.cn
@@ 1 Introduction
"No health without mental health" is becoming more than a slogan, with approximately 14% of the global disease burden attributed to neuropsychiatric disorders Prince et al. (2007). Despite the affordability and effectiveness of many mental health treatments, a significant gap persists between those in need and those able to access care Freeman (2022). The World Health Organization (WHO) continually advocates for increased investment to augment understanding and dispel stigma associated with mental health disorders. Yet, the challenge of ensuring quality, affordable care for mental health conditions remains formidable. Consequently, the identification of novel treatments and enhancement of existing therapies for all mental diseases are key objectives in the research domain.
@@ 3.2 Data Processing
Considering the differences in data sources of our collection, we need to reformat these collected reports according to a uniform standard. To build a comprehensive and helpful dataset, we combine the case format of China’s National Class 2 Psychological Counselor Examination and other psychological counseling literature to regularize collected reports, where the following 6 components are included: Title, Type, Method, Case Brief, Consultation Process and Experience Thoughts. Note that the consultation process is written from a third-person perspective and does not contain specific dialog. For detailed description of components and examples of psychological counseling report, please refer to Appendix A.
@@ 4.2 Intrinsic Evaluation of CPsyCoun
To ensure comprehensiveness and diversity of the evaluation dataset, we randomly select 20 cases from each of the seven classic schools of psychological counseling in CPsyCounR, acquiring a total of 140 cases. Then we adopt direct role-play prompting and Memo2Demo method respectively for dialogue generation, and instruct GPT-4 Achiam et al. (2023) to conduct a comparative evaluation of the above two multi-turn consultation dialogues. The evaluation standard refers to the evaluation metrics shows in Table 4. For detailed evaluation prompt, please refer to Figure 9 in the appendix.
@@ 4.3 Extrinsic Evaluation of CPsyCoun
Figure 4: Radar plot of detailed scores of CPsyCounX and other baselines on 9 counseling topics on the benchmark.
@@
The turn-based dialogue evaluation method is adopted on CPsyCounE for the following extrinsic evaluation. We include InternLM2-7B-Chat Team (2023), SoulChat Chen et al. (2023b), ChatGPT and GLM-4 Zeng et al. (2023) as major baseline models. The evaluation standard refers to the evaluation metrics in Table 4 in the appendix. To accommodate multi-turn dialogues, we adjust the authenticity criterion accordingly. For detailed evaluation prompt, please refer to Figure 10 in the appendix.
@@ Acknowledgements
This work was partially supported by the China Postdoctoral Science Foundation (2023M733654) and Guangdong Basic and Applied Basic Research Foundation (2023A1515110496).
@@ Appendix A Unified Format of Psychological Counseling Reports
Considering the differences in data sources of our collection, we need to reformat these collected reports according to a uniform standard. To regularize the psychological counseling reports collected from different data sources, we include six components within each report. In Figure 5, we offer detailed description of these components, accompanied by real examples from collected reports. Note that the original case is written in Chinese, while the English version is translated by us for reference.

### v3

@@ Abstract
Chenhao Zhang1,3thanks: Equal contribution.thanks: Work done on the Science and Technology Innovation Project of UCAS directed by SIAT.and Renhao Li2,311footnotemark: 1and Minghuan Tan3thanks: Corresponding author.and Min Yang [^3] footnotemark: 3and Jingwei Zhu4and Di Yang4and Jiahao Zhao3,522footnotemark: 2and Guancheng Ye622footnotemark: 2and Chengming Li7and Xiping Hu7 1 Huazhong University of Science and Technology 2 University of Macau 3 Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 4 University of Science and Technology of China 5 Jilin University 6 South China University of Technology 7 Shenzhen MSU-BIT University ch_zhang@hust.edu.cn, li.renhao@connect.um.edu.mo {mh.tan,min.yang}@siat.ac.cn
@@ 1 Introduction
"No health without mental health" is becoming more than a slogan, with approximately 14% of the global disease burden attributed to neuropsychiatric disorders Prince et al. (2007). Despite the affordability and effectiveness of many mental health treatments, a significant gap persists between those in need and those able to access care Freeman (2022). The World Health Organization (WHO) continually advocates for increased investment to augment understanding and dispel the stigma associated with mental health disorders. Yet, the challenge of ensuring quality, affordable care for mental health conditions remains formidable. Consequently, the identification of novel treatments and enhancement of existing therapies for all mental diseases are key objectives in the research domain.
@@ 3.2 Data Processing
Considering the differences in data sources of our collection, we need to reformat these collected reports according to a uniform standard. To build a comprehensive and helpful dataset, we combine the case format of China’s National Class 2 Psychological Counselor Examination and other psychological counseling literature to regularize collected reports, where the following 6 components are included: Title, Type, Method, Case Brief, Consultation Process and Experience Thoughts. Note that the consultation process is written from a third-person perspective and does not contain specific dialog. For a detailed description of components and examples of psychological counseling report, please refer to Appendix A.
@@ 4.2 Intrinsic Evaluation of CPsyCoun
To ensure comprehensiveness and diversity of the evaluation dataset, we randomly select 20 cases from each of the seven classic schools of psychological counseling in CPsyCounR, acquiring a total of 140 cases. Then we adopt direct role-play prompting and Memo2Demo method respectively for dialogue generation, and instruct GPT-4 Achiam et al. (2023) to conduct a comparative evaluation of the above two multi-turn consultation dialogues. The evaluation standard refers to the evaluation metrics shows in Table 4. For detailed evaluation prompts, please refer to Figure 9 in the appendix.
@@ 4.3 Extrinsic Evaluation of CPsyCoun
Figure 4: Radar plot of detailed scores of CPsyCounX and other baselines on 9 counseling topics on CPsyCounE.
@@
The turn-based dialogue evaluation method is adopted on CPsyCounE for the following extrinsic evaluation. We include InternLM2-7B-Chat Team (2023), SoulChat Chen et al. (2023b), ChatGPT and GLM-4 Zeng et al. (2023) as major baseline models. The evaluation standard refers to the evaluation metrics in Table 4 in the appendix. To accommodate multi-turn dialogues, we adjust the authenticity criterion accordingly. For detailed evaluation prompts, please refer to Figure 10 in the appendix.
@@ Acknowledgements
This work was partially supported by National Key Research and Development Program of China (2022YFF0902100), China Postdoctoral Science Foundation (2023M733654), National Natural Science Foundation of China (62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166), Shenzhen Science and Technology Innovation Program (KQTD20190929172835662), Shenzhen Basic Research Foundation (JCYJ20210324115614039).
@@ Appendix A Unified Format of Psychological Counseling Reports
Considering the differences in data sources of our collection, we need to reformat these collected reports according to a uniform standard. To regularize the psychological counseling reports collected from different data sources, we include six components within each report. In Figure 5, we offer a detailed description of these components, accompanied by real examples from collected reports. Note that the original case is written in Chinese, while the English version is translated by us for reference.